ABC Company is a discount department store that has three major departments, groceries, general   merchandise and drugs. Management is considering dropping groceries, which have consistently shown   a net loss. The space vacated by the dropping of groceries would remain idle and the common fixed   cost would remain the same. The following is the financials for the three departments: Total Groceries Merchandise Drugs Sales $       1,900 $         1,000 $               800 $            100 Variable costs $       1,420 $            800 $               560 $              60 Contribution margin $          480 $            200 $               240 $              40 Fixed expenses Direct fixed costs $          265 $            150 $               100 $              15 Common fixed costs $          180 $              60 $               100 $              20 Operating income $            35 $                     (10) $                 40 $                5 Required: a) Compute the change of operating income if the company drop groceries. b) Should the company drop groceries?    c) Assume the company wants to use the space made available by the dropping of groceries to expand the general merchandise department. The general merchandise expansion would increase sales and costs as follows: Increase in sales   $          500 Increase in direct fixed costs $            70 The contribution margin of the increased general merchandise sales remains the same as before the expansion. Required: Compute the change in operating income if the company decides to drop the groceries department and expand the general merchandise department.
Answer #1

Change in Operating Income by dropping groceries = Fixed costs avoided – Contribution Margin lost

= Direct Fixed costs – Contribution Margin lost

= 150-200

= -$50

Hence, Operating Income would decrease by $50

b)No, should not drop as it will lead to a fall in net income

c) Contribution Margin on Merchandise Sales = 240/800 = 30%

Change in Operating Income = (Increase in contribution margin from general merchandise sales) - Increase in Direct fixed costs - Loss on dropping groceries = 500*30% - 70 - 50

= $30 Increase